# Towards Robust and Efficient Cloud-Edge Elastic Model Adaptation via Selective Entropy Distillation

### v2

@@ 3 Experiments
Implementation details. We set the entropy thresholds as the initialized value following (Niu et al., 2022) and , where denotes the number of classes. Then the threshold decreases based on Eqn. (2) with . For the adaptation of the foundation and edge model, we use both an SGD optimizer with a learning rate of and a momentum of . For the adaptation of the edge model, we set the batch size to , in which samples are newly uploaded and the remaining samples are randomly sampled from the replay buffer. The hyper-parameter and are both set to . We will release our code publicly upon acceptance.
@@
| BN Adaptation | 42.3 | 39.8 | 40.0 | 37.5 | 31.4 | 45.1 | 44.3 | 40.8 | 36.2 | 53.9 | 65.0 | 58.2 | 60.2 | 58.0 | 57.7 | 47.4 |
| ONDA | 40.0 | 38.9 | 37.5 | 29.5 | 27.5 | 43.8 | 43.9 | 40.2 | 35.2 | 54.6 | 65.1 | 56.1 | 59.7 | 58.6 | 57.6 | 45.9 |
| LAME | 20.6 | 18.9 | 17.2 | 29.5 | 14.7 | 28.3 | 26.9 | 26.8 | 23.2 | 34.9 | 62.4 | 37.5 | 51.3 | 41.1 | 52.5 | 32.4 |
@@
| BN Adaptation | 16.6 | 16.2 | 17.3 | 18.6 | 18.2 | 25.9 | 34.7 | 28.4 | 29.8 | 41.2 | 58.5 | 22.2 | 40.1 | 45.3 | 38.0 | 30.1 |
| ONDA | 13.7 | 15.0 | 14.1 | 12.3 | 13.2 | 23.7 | 34.2 | 29.4 | 28.6 | 40.9 | 58.5 | 12.3 | 39.3 | 44.6 | 37.5 | 27.8 |
| LAME | 0.9 | 1.1 | 0.6 | 11.2 | 8.2 | 10.8 | 17.0 | 8.7 | 15.6 | 12.4 | 51.1 | 3.3 | 14.9 | 22.5 | 30.1 | 13.9 |
@@ 3.2 Performance Comparisons on ImageNet-R
| BN Adapt. | 22.8 | – |
| ONDA | 22.7 | – |
| LAME | 20.2 | – |
@@ C.2 More experimental protocols
For the adaptation of the edge model, we set the batch size to , in which samples are newly uploaded and the remaining samples are randomly sampled from the replay buffer. We set the hyper-parameters and is set to and , respectively. We set the replay buffer size to 10,000. Note that all the models used in our experiments are pretrained on ImageNet training set and available publicly. Specifically, ResNet18, ResNet101, MobileNetV2, MobileNetV3 and ShuffleNetV2 are from torchvision.333https://github.com/pytorch/vision DeiT-tiny and DeiT-base are from facebookresearch/deit.444https://github.com/facebookresearch/deit CLIP-ViT-B/32 is from openai/CLIP.555https://github.com/openai/CLIP Since the edge model may infer with a small batch size (e.g., the batch size is 1), we introduce Batch Renormalization (Ioffe, 2017) to replace the vanilla Batch Normalization in the edge model following (Zhao et al., 2023). In this case, is able to infer over a very small batch of test samples with moving average statistics instead of batch re-computing statistics. We will release our code publicly upon acceptance.
@@ D.1 More comparisons with CNN-based models on ImageNet-C
| BN Adaptation | 59.3 | 57.9 | 51.9 | 57.2 | 57.4 | 62.4 | 54.9 | 54.1 | 57.7 | 61.3 | 67.9 | 64.9 | 62.2 | 64.8 | 63.2 | 59.8 |
| ONDA | 58.7 | 58.0 | 51.3 | 55.2 | 56.1 | 62.3 | 55.0 | 53.6 | 57.2 | 61.9 | 68.3 | 64.6 | 62.3 | 65.1 | 63.0 | 59.5 |
| LAME | 48.8 | 46.3 | 34.1 | 51.4 | 48.2 | 56.8 | 43.6 | 45.9 | 52.3 | 52.5 | 67.3 | 57.8 | 60.1 | 59.5 | 58.9 | 52.2 |
@@
| BN Adaptation | 52.5 | 49.6 | 45.5 | 50.4 | 48.7 | 55.7 | 48.6 | 39.4 | 45.1 | 58.3 | 66.8 | 62.7 | 46.2 | 63.8 | 60.2 | 52.9 |
| ONDA | 51.3 | 49.1 | 43.6 | 45.5 | 45.7 | 55.2 | 48.4 | 39.9 | 44.1 | 59.0 | 67.2 | 61.9 | 46.5 | 63.4 | 60.1 | 52.1 |
| LAME | 37.2 | 33.0 | 24.7 | 43.7 | 35.8 | 44.9 | 33.6 | 22.6 | 34.5 | 45.4 | 65.4 | 50.8 | 38.8 | 59.2 | 55.3 | 41.7 |
@@
| BN Adaptation | 29.8 | 24.9 | 28.2 | 26.8 | 25.7 | 33.0 | 39.2 | 31.4 | 35.0 | 50.5 | 62.2 | 44.6 | 54.4 | 49.0 | 49.0 | 38.9 |
| ONDA | 26.9 | 23.6 | 24.8 | 19.0 | 21.4 | 30.7 | 38.7 | 31.2 | 33.8 | 51.4 | 62.3 | 37.4 | 54.0 | 49.5 | 48.7 | 36.9 |
| LAME | 6.9 | 5.2 | 5.1 | 18.3 | 11.4 | 15.8 | 21.4 | 16.8 | 21.3 | 27.3 | 57.6 | 13.4 | 38.6 | 25.6 | 43.1 | 21.8 |

### v3

@@ 3 Experiments
Implementation details. We set the entropy thresholds as the initialized value following (Niu et al., 2022) and , where denotes the number of classes. Then the threshold decreases based on Eqn. (2) with . For the adaptation of the foundation and edge model, we use both an SGD optimizer with a learning rate of and a momentum of . For the adaptation of the edge model, we set the batch size to , in which samples are newly uploaded and the remaining samples are randomly sampled from the replay buffer. The hyper-parameter and are both set to .
@@
| BN Adaptation† | 42.3 | 39.8 | 40.0 | 37.5 | 31.4 | 45.1 | 44.3 | 40.8 | 36.2 | 53.9 | 65.0 | 58.2 | 60.2 | 58.0 | 57.7 | 47.4 |
| ONDA† | 40.0 | 38.9 | 37.5 | 29.5 | 27.5 | 43.8 | 43.9 | 40.2 | 35.2 | 54.6 | 65.1 | 56.1 | 59.7 | 58.6 | 57.6 | 45.9 |
| LAME† | 20.6 | 18.9 | 17.2 | 29.5 | 14.7 | 28.3 | 26.9 | 26.8 | 23.2 | 34.9 | 62.4 | 37.5 | 51.3 | 41.1 | 52.5 | 32.4 |
@@
| BN Adaptation† | 16.6 | 16.2 | 17.3 | 18.6 | 18.2 | 25.9 | 34.7 | 28.4 | 29.8 | 41.2 | 58.5 | 22.2 | 40.1 | 45.3 | 38.0 | 30.1 |
| ONDA† | 13.7 | 15.0 | 14.1 | 12.3 | 13.2 | 23.7 | 34.2 | 29.4 | 28.6 | 40.9 | 58.5 | 12.3 | 39.3 | 44.6 | 37.5 | 27.8 |
| LAME† | 0.9 | 1.1 | 0.6 | 11.2 | 8.2 | 10.8 | 17.0 | 8.7 | 15.6 | 12.4 | 51.1 | 3.3 | 14.9 | 22.5 | 30.1 | 13.9 |
@@ 3.2 Performance Comparisons on ImageNet-R
| BN Adapt.† | 22.8 | – |
| ONDA† | 22.7 | – |
| LAME† | 20.2 | – |
@@
| Model | Accuracy (%) | #Uploaded samples |
| ResNet18 (baseline) | 33.1 | – |
| Tent | 30.5 | 750k (100%) |
| ETA | 44.5 | 384k (51%) |
| CEMA (Ours) | 45.6 | 286k (38%) |
| Method | Level 3 | Level 5 | ||
| Acc. (%) | #Upload | Acc. (%) | #Upload | |
| w/ Fixed | 51.2 | 25,325 | 30.0 | 15,473 |
| Dynamic | 51.1 | 20,976 | 29.9 | 10,081 |
| 51.1 | 17,479 | 29.8 | 9,889 | |
| Replay Buffer | Level 3 | Level 5 | ||
| Acc. (%) | #Upload | Acc. (%) | #Upload | |
| 47.4 | 17,654 | 25.0 | 9,084 | |
| 51.1 | 17,479 | 29.8 | 9,889 | |
@@ C.2 More experimental protocols
For the adaptation of the edge model, we set the batch size to , in which samples are newly uploaded and the remaining samples are randomly sampled from the replay buffer. We set the hyper-parameters and is set to and , respectively. We set the replay buffer size to 10,000. Note that all the models used in our experiments are pretrained on ImageNet training set and available publicly. Specifically, ResNet18, ResNet101, MobileNetV2, MobileNetV3 and ShuffleNetV2 are from torchvision.333https://github.com/pytorch/vision DeiT-tiny and DeiT-base are from facebookresearch/deit.444https://github.com/facebookresearch/deit CLIP-ViT-B/32 is from openai/CLIP.555https://github.com/openai/CLIP Since the edge model may infer with a small batch size (e.g., the batch size is 1), we introduce Batch Renormalization (Ioffe, 2017) to replace the vanilla Batch Normalization in the edge model following (Zhao et al., 2023). In this case, is able to infer over a very small batch of test samples with moving average statistics instead of batch re-computing statistics.
@@ D.1 More comparisons with CNN-based models on ImageNet-C
| BN Adaptation† | 59.3 | 57.9 | 51.9 | 57.2 | 57.4 | 62.4 | 54.9 | 54.1 | 57.7 | 61.3 | 67.9 | 64.9 | 62.2 | 64.8 | 63.2 | 59.8 |
| ONDA† | 58.7 | 58.0 | 51.3 | 55.2 | 56.1 | 62.3 | 55.0 | 53.6 | 57.2 | 61.9 | 68.3 | 64.6 | 62.3 | 65.1 | 63.0 | 59.5 |
| LAME† | 48.8 | 46.3 | 34.1 | 51.4 | 48.2 | 56.8 | 43.6 | 45.9 | 52.3 | 52.5 | 67.3 | 57.8 | 60.1 | 59.5 | 58.9 | 52.2 |
@@
| BN Adaptation† | 52.5 | 49.6 | 45.5 | 50.4 | 48.7 | 55.7 | 48.6 | 39.4 | 45.1 | 58.3 | 66.8 | 62.7 | 46.2 | 63.8 | 60.2 | 52.9 |
| ONDA† | 51.3 | 49.1 | 43.6 | 45.5 | 45.7 | 55.2 | 48.4 | 39.9 | 44.1 | 59.0 | 67.2 | 61.9 | 46.5 | 63.4 | 60.1 | 52.1 |
| LAME† | 37.2 | 33.0 | 24.7 | 43.7 | 35.8 | 44.9 | 33.6 | 22.6 | 34.5 | 45.4 | 65.4 | 50.8 | 38.8 | 59.2 | 55.3 | 41.7 |
@@
| BN Adaptation† | 29.8 | 24.9 | 28.2 | 26.8 | 25.7 | 33.0 | 39.2 | 31.4 | 35.0 | 50.5 | 62.2 | 44.6 | 54.4 | 49.0 | 49.0 | 38.9 |
| ONDA† | 26.9 | 23.6 | 24.8 | 19.0 | 21.4 | 30.7 | 38.7 | 31.2 | 33.8 | 51.4 | 62.3 | 37.4 | 54.0 | 49.5 | 48.7 | 36.9 |
| LAME† | 6.9 | 5.2 | 5.1 | 18.3 | 11.4 | 15.8 | 21.4 | 16.8 | 21.3 | 27.3 | 57.6 | 13.4 | 38.6 | 25.6 | 43.1 | 21.8 |